# Neurodevelopmental disorders modeling using isogeometric analysis, dynamic domain expansion and local refinement

### v1

@@ Abstract
Neurodevelopmental disorders (NDDs) have arisen as one of the most prevailing chronic diseases within the US. Often associated with severe adverse impacts on the formation of vital central and peripheral nervous systems during the neurodevelopmental process, NDDs are comprised of a broad spectrum of disorders, such as autism spectrum disorder, attention deficit hyperactivity disorder, and epilepsy, characterized by progressive and pervasive detriments to cognitive, speech, memory, motor, and other neurological functions in patients. However, the heterogeneous nature of NDDs poses a significant roadblock to identifying the exact pathogenesis, impeding accurate diagnosis and the development of targeted treatment planning. A computational NDDs model holds immense potential in enhancing our understanding of the multifaceted factors involved and could assist in identifying the root causes to expedite treatment development. To tackle this challenge, we introduce optimal neurotrophin concentration to the driving force and degradation of neurotrophin to the synaptogenesis process of a 2D phase field neuron growth model using isogeometric analysis to simulate neurite retraction and atrophy. The optimal neurotrophin concentration effectively captures the inverse relationship between neurotrophin levels and neurite survival, while its degradation regulates concentration levels. Leveraging dynamic domain expansion, the model efficiently expands the domain based on outgrowth patterns to minimize degrees of freedom. Based on truncated T-splines, our model simulates the evolving process of complex neurite structures by applying local refinement adaptively to the cell/neurite boundary. Furthermore, a thorough parameter investigation is conducted with detailed comparisons against neuron cell cultures in experiments, enhancing our fundamental understanding of the mechanisms underlying NDDs.
@@ 1 Introduction
Neurodevelopmental disorders (NDDs), often associated with impairments during the neuron developmental process, pose persistent complications that have a profound and long-lasting impact on patients [1, 2]. A wide range of potential factors behind the etiology of NDDs are gaining recognition recently, with emerging research highlighting their significant impact on the pathophysiology of these conditions [3, 4]. Investigations into the functional roles of these factors have revealed some potential influences on NDDs progression, indicating that they play crucial roles during the neurodevelopmental process [5, 6]. Moreover, emerging evidence suggests that certain factors involved during the neurodevelopmental process may play protective roles, potentially preventing the onset of NDDs [7, 8]. However, the lack of comprehensive disorder studies of the specific biological functions or biophysical processes highlights a significant gap in the current field. Furthermore, NDDs often manifest through complex neuron morphological transformations, such as beading, retraction, and atrophy. All of these behaviors add complexity to a thorough neurological study and pose challenges to therapeutic targeting and planning [9, 10]. Analyzing the complex morphological changes observed in NDDs and understanding the biological mechanisms that drive these conditions are crucial for developing effective therapeutic strategies. Given the level of complexity associated with the sophisticated processes involved, unconventional approaches are necessary to enhance our understanding of the NDDs process. As such, there is a pressing need for a comprehensive computational model to study NDDs and unravel the intricate neurodevelopmental process. These models will be indispensable for shedding light on the fundamental causes and complexities of NDDs, consequently enabling the development of precise and potent treatments [11, 12].
@@
In computational neuroscience, significant progress has been made in mathematically modeling neurodevelopmental processes, including initial neurite outgrowth [13], axon differentiation [14, 15] and axon guidance [16]. Phenomenological models have provided insights into a wide range of neurodevelopmental processes such as filopodia [17], external repulsive cues [18], stochastic mechanisms [19], generalized neurite characteristics for modeling morphology [20, 21], and incorporating interactions with surrounding substrates [22]. However, these models focus on phenomenological outcomes, often overlooking underlying biophysics [23, 24]. Integrating comprehensive biophysical mechanisms into neuron growth models comes with significant computational costs and numerical instabilities [25, 26], which are particularly aggravated in scenarios like intracellular material transport [27] and traffic jams in complex 3D structures [28]. Despite these complications, ongoing efforts are underway to enhance and validate computational models with biophysical phenomena from experimental observations. Among different approaches, some noteworthy work includes the utilization of phase field techniques for neuron growth modeling [29], exploration of neurotrophin interactions [30], and the development of biophysically coupled phase field neuron growth models [31]. There are also models that integrate neurite morphometric features and bridge the gap between theoretical and experimental neuroscience [32, 33]. Considering the complexity and large variety of the neural development process [34], robust computational approaches that tackle high-order equations on complex geometries are crucial for advancing our understanding of neurodevelopmental processes and paving the way toward effective targeted therapeutic interventions. This raises the need for isogeometric analysis (IGA) with high-fidelity spline modeling techniques that provide the necessary flexibility and precision [35, 36], particularly through domain expansion and localized refinements to accurately capture the dynamic and detailed evolution of neuron growth during development and degradation during apoptosis on response to damage or pathology.
@@
    Introduction of optimal neurotrophin concentration into the driving force and degradation of neurotrophin into the synaptogenesis process of the phase field model to simulate neurite retraction and atrophy. The optimal neurotrophin concentration effectively captures the inverse relationship between neurotrophin levels and neurite survival, and its degradation regulates concentration levels, providing critical insights into NDDs;
@@
    Comprehensive NDDs study utilizing the computational model to investigate neurite retraction and atrophy, providing new insights into possible factors and biophysical mechanisms underlying NDDs; and
@@
    Detailed validation through comparisons of the external-cue guided computational NDDs model with experimental healthy human induced pluripotent stem cell (iPSC)-derived neurons and rat hippocampus neurons undergoing excitotoxic apoptosis.
@@ 3 IGA-based phase field NDDs modeling
    Driving force equation that couples the effect of tubulin and neurotrophin back to the phase field equation. We introduce optimal neurotrophin concentration to model the inverse relationship between neurotrophin level and neurite survival. This inverse relationship adjusts the driving force magnitude and, therefore, drives the phase field interface evolution for NDDs modeling.
@@
$\frac{\partial\phi}{\partial t}=M_{\phi}\left[\bigtriangledown\cdot\left(a(\Psi)^{2}\bigtriangledown\phi\right)+4W\phi(1-\phi)\left(\phi-\frac{1}{2}+\frac{15}{2W}\phi(1-\phi)(f_{L}-f_{S})\right)\right].$ (7)
@@
$\frac{\partial\phi}{\partial t}=M_{\phi}\left[\bigtriangledown\cdot\left(a(\Psi)^{2}\bigtriangledown\phi\right)-\frac{\partial}{\partial{x}}\left(a(\Psi)\frac{\partial a(\Psi)}{\partial\Psi}\frac{\partial\phi}{\partial y}\right)+\frac{\partial}{\partial{y}}\left(a(\Psi)\frac{\partial a(\Psi)}{\partial\Psi}\frac{\partial\phi}{\partial x}\right)+\phi(1-\phi)\left(\phi-\frac{1}{2}+F_{driv}+6H|\bigtriangledown\theta|\right)\right],$ (9)
@@
$\frac{\partial c_{neur}}{\partial t}=D_{c}\bigtriangledown^{2}c_{neur}+(K-k_{p75}c_{neur})\frac{\partial\phi}{\partial t}-k_{2}c_{neur},$ (12)
@@
Driving force equation. In neurons, neurotrophin concentrations are a determining factor in neuronal survival. An optimal amount of neurotrophin is necessary for the survival of neurons [86]. Lower levels typically support neuron survival, while higher concentrations can negate this effect. This inverse relationship between neurotrophin concentrations and cell survival is particularly notable at higher concentrations, where an observed survival-promoting effect at lower concentrations is reversed [87]. To model the aforementioned inverse relationship between neurotrophin concentrations and cell survival considering the intricate balance of neurotrophin concentration [88], we introduce $c_{opti}$ to the driving force equation and obtain
@@
where $\frac{\alpha}{\pi}$ is a scaling coefficient, $H_{\epsilon}$ is a Heaviside step function, and $\gamma$ is the interfacial energy constant. By introducing $c_{opti}$, Eqn. 13 can effectively model the inverse relationship between $c_{neur}$ and neurite survival. When the magnitude of $c_{neur}$ exceeds $c_{opti}$, the effect of $c_{neur}$ is reversed, leading to the $\phi$ interface retractions.
@@ 4.1 Review of truncated T-spline
where $c_{i}$ as the refinement coefficients from knot insertion [89], and $N_{U^{\prime}_{i}}$ are the active children of $N_{U^{\prime}}(u)$. A bi-variate T-spline basis function, $B(u,v)=N_{U}(u)N_{V}(v)$, is the tensor product of two uni-variate functions, and its refinement is achieved by refining $N_{U}(u)$ and $N_{V}(v)$ [63].
@@
where $P_{ij}$ are the control points and $trunB_{ij}(u,v)$ are the truncated T-spline basis functions. Note that $trunB_{ij}(u,v)$ at any particular point sums up to 1 to ensure the property of partition of unity [93], which makes truncated T-splines particularly effective for precise computational analysis with complex geometries. We refer readers to [63] for an in-depth discussion on truncated T-splines.
@@ 4.2 $\phi$ interface-based local refinements
Figure 2: Local refinement using truncated T-splines. (A) Phase field variable $\phi$ initialization on a coarse mesh. (B) Identifying elements at the $\phi$ interface. (C) Interface elements are locally refined with subdivision. (D) Detecting face-face intersections created by face extension from subdivisions. (E) Applying 1-ring of bisections to resolve these face-face intersections and maintaining the integrity of the mesh for truncated T-splines. (F) Generating locally refined T-mesh for truncated T-splines with enhanced precision in regions of interest. (G) Interpolating $\phi$ initialization from coarse control mesh to the locally-refined mesh. (H) The locally refined $\phi$ initialization.
To solve the phase field model, $\phi$ initialization is needed. To apply local refinements, the model first needs to know $\phi$ initialization on a uniform coarse control mesh (Fig. 2A), refine elements at $\phi$ interface, and then obtain $\phi$ initialization on the locally refined mesh. Beginning with a uniform coarse control mesh, we can calculate $\phi$ initialization based on the initial cell center and cell radius $r_{0}$. Then, by selecting elements along $\phi$ interface (Fig. 2B), we can refine and enhance the effective mesh resolution at these critical locations of interest (Fig. 2C) while keeping the mesh coarse in areas where $\phi$ values are stable at either 0 or 1. This targeted refinement approach ensures efficient use of computational resources. Because a valid T-mesh in truncated T-splines must be strongly balanced and free of face-face intersections to satisfy linear independence [63, 90], we need to resolve these face-face intersections in the automated local refinement process (Fig. 2D). To address this issue, we apply bisections to 1-ring of elements surrounding the selected interface elements (Fig. 2E). This approach effectively resolves face-face intersections and allows the model to refine the mesh only where needed. Thus, we can obtain a locally-refined T-mesh for truncated T-splines tailored explicitly to the given $\phi$ value (Fig. 2F). Subsequently, $\phi$ values are interpolated and transferred onto this newly refined mesh using the KD tree-based interpolation method [94], providing accurate initialization for the computation of the evolving neuron outgrowth $\phi$ (Fig. 2G). Starting with an existing control point, a search using the KD tree data structure finds the nearest new control point. If a matching control point is found, the value at the input control point is directly used. If the new control point is situated among the existing points, the algorithm interpolates depending on their positions and connectivity in the control mesh (Fig. 2G). This conditional handling ensures values are correctly passed to the locally refined mesh. The KD tree approach effectively utilizes spatial indexing and positions, preserving the initialization accuracy while minimizing unnecessary computations.
@@ 5 Dynamic domain expansion
The model incorporates a directional domain expansion algorithm (Algorithm 1) to dynamically adjust the control mesh in response to neurite approaching the domain boundary (double-sided arrows in Fig. 3A-D). The algorithm takes neuron growth $\phi$ and control mesh as the input and evaluates each element adjacent to the domain boundary, searching for any nonzero values of the $\phi$. If nonzero $\phi$ is detected near the boundary, the algorithm flags the corresponding edge for expansion using $expFlag$. Once boundaries are flagged, the algorithm modifies the domain size based on the predetermined expansion size (default $3\times\Delta x$, where $\Delta x$ is coarse element size). It generates an enlarged control mesh (dashed zone in Figure 3A-D) and translates the entire control mesh according to the specified direction of expansion to maintain the position of the neuron $\phi$. This directional domain expansion approach works in tandem with the truncated T-splines generation, discussed in Section 4, to ensure that the computational domain adapts to accommodate growing neurites with the least amount of elements, maintaining analysis accuracy and efficiency. The transition to the expanded mesh requires an interpolation operation that passes values from the original control points to the new control points (Fig. 2G). The interpolation process for a dynamically expanding mesh in the NDDs simulation leverages the KD tree algorithm to optimize spatial queries of control points $cpts$, significantly accelerating the search for the nearest $cpts$ [94]. Starting with new control points $cpts^{\prime}$, each $cpts^{\prime}$ undergoes a rounding process to identify whether it is located on expanded elements. This procedure significantly speeds up the algorithm by skipping elements that are not expanded, and reduces the total number of DOFs while preserving the analysis accuracy.

### v2

@@ Abstract
Neurodevelopmental disorders (NDDs) have arisen as one of the most prevailing chronic diseases within the US. Often associated with severe adverse impacts on the formation of vital central and peripheral nervous systems during the neurodevelopmental process, NDDs are comprised of a broad spectrum of disorders, such as autism spectrum disorder, attention deficit hyperactivity disorder, and epilepsy, characterized by progressive and pervasive detriments to cognitive, speech, memory, motor, and other neurological functions in patients. However, the heterogeneous nature of NDDs poses a significant roadblock to identifying the exact pathogenesis, impeding accurate diagnosis and the development of targeted treatment planning. A computational NDDs model holds immense potential in enhancing our understanding of the multifaceted factors involved and could assist in identifying the root causes to expedite treatment development. To tackle this challenge, we introduce optimal neurotrophin concentration to the driving force and degradation of neurotrophin to the synaptogenesis process of a 2D phase field neuron growth model using isogeometric analysis to simulate neurite retraction and atrophy. The optimal neurotrophin concentration effectively captures the inverse relationship between neurotrophin levels and neuron survival, while its degradation regulates concentration levels. Leveraging dynamic domain expansion, the model efficiently expands the domain based on outgrowth patterns to minimize degrees of freedom. Based on truncated T-splines, our model simulates the evolving process of complex neurite structures by applying local refinement adaptively to the cell/neurite boundary. Furthermore, a thorough parameter investigation is conducted with detailed comparisons against neuron cell cultures in experiments, enhancing our fundamental understanding of the possible mechanisms underlying NDDs.
@@ 1 Introduction
Neurodevelopmental disorders (NDDs), often associated with impairments during the neuron developmental process, pose persistent complications that have a profound and long-lasting impact on patients such as attention deficit hyperactivity disorder (ADHD) and autism spectrum disorder [1, 2]. A wide range of potential factors behind the etiology of NDDs are gaining recognition recently, with emerging research highlighting their significant impact on the pathophysiology of these conditions [3, 4]. Investigations into the functional roles of these factors have revealed some potential influences on NDDs progression, indicating that they play crucial roles during the neurodevelopmental process [5, 6]. Moreover, emerging evidence suggests that certain factors involved during the neurodevelopmental process may play protective roles, potentially preventing the onset of NDDs [7, 8]. However, the lack of comprehensive disorder studies of the specific biological functions or biophysical processes highlights a significant gap in the current field. Furthermore, NDDs often manifest through complex neuron morphological transformations, such as beading, retraction, and atrophy. All of these behaviors add complexity to a thorough neurological study and pose challenges to therapeutic targeting and planning [9, 10]. Analyzing the complex morphological changes observed in NDDs and understanding the biological mechanisms that drive these conditions are crucial for developing effective therapeutic strategies. Given the level of complexity associated with the sophisticated processes involved, unconventional approaches are necessary to enhance our understanding of the NDDs process. As such, there is a pressing need for a comprehensive computational model to study NDDs and unravel the intricate neurodevelopmental process. These models will be indispensable for shedding light on the fundamental causes and complexities of NDDs, consequently enabling the development of precise and potent treatments [11, 12].
@@
In computational neuroscience, significant progress has been made in mathematically modeling neurodevelopmental processes, including initial neurite outgrowth [13], axon differentiation [14, 15] and axon guidance [16]. Phenomenological models have provided insights into a wide range of neurodevelopmental processes such as filopodia [17], external repulsive cues [18], stochastic mechanisms [19], generalized neurite characteristics for modeling morphology [20, 21], and incorporating interactions with surrounding substrates [22]. However, these models focus on phenomenological outcomes, often overlooking underlying biophysics [23, 24]. Integrating comprehensive biophysical mechanisms into neuron growth models comes with significant computational costs and numerical instabilities [25, 26], which are particularly aggravated in scenarios like intracellular material transport [27] and traffic jams in complex 3D structures [28]. Despite these complications, ongoing efforts are underway to enhance and validate computational models with biophysical phenomena from experimental observations. Among different approaches, some noteworthy work includes the utilization of phase field techniques for neuron growth modeling [29], exploration of neurotrophin interactions [30], and the development of biophysically coupled phase field neuron growth models [31]. There are also models that integrate neurite morphometric features and bridge the gap between theoretical and experimental neuroscience [32, 33]. Considering the complexity and large variety of the neurodevelopmental process [34], robust computational approaches that tackle high-order equations on complex geometries are crucial for advancing our understanding of neurodevelopmental processes and paving the way toward effective targeted therapeutic interventions. This raises the need for isogeometric analysis (IGA) with high-fidelity spline modeling techniques that provide the necessary flexibility and precision [35, 36], particularly through domain expansion and localized refinements to accurately capture the dynamic and detailed evolution of neuron growth during development and degeneration in response to damage or pathology.
@@
    Introduction of optimal neurotrophin concentration into the driving force and degradation of neurotrophin into the synaptogenesis process of the phase field model to simulate neurite retraction and atrophy. The optimal neurotrophin concentration effectively captures the inverse relationship between neurotrophin levels and neuron survival, and its degradation regulates concentration levels, providing hypotheses into NDD mechanisms;
@@
    Comprehensive NDDs model parameter study utilizing the computational model to investigate neurite retraction and atrophy, providing new insights into possible factors and biophysical mechanisms underlying NDDs; and
@@
    Detailed qualitative validation through comparisons of the external-cue guided computational NDDs model with experimental healthy human induced pluripotent stem cell (iPSC)-derived neurons and rat hippocampus neurons undergoing cell death and neurite fragmentation.
@@ 3 IGA-based phase field NDDs modeling
    Driving force equation that couples the effect of tubulin and neurotrophin back to the phase field equation. We introduce optimal neurotrophin concentration to model the inverse relationship between neurotrophin level and neuron survival. This inverse relationship adjusts the driving force magnitude and, therefore, drives the phase field interface evolution for NDDs modeling.
@@
$\frac{\partial\phi}{\partial t}=M_{\phi}\left[\bigtriangledown\cdot\left(a(\Psi)^{2}\nabla\phi\right)+4W\phi(1-\phi)\left(\phi-\frac{1}{2}+\frac{15}{2W}\phi(1-\phi)(f_{L}-f_{S})\right)\right].$ (7)
@@
$\frac{\partial\phi}{\partial t}=M_{\phi}\left[\nabla\cdot\left(a(\Psi)^{2}\nabla\phi\right)-\frac{\partial}{\partial x}\left(a(\Psi)\frac{\partial a(\Psi)}{\partial\Psi}\frac{\partial\phi}{\partial y}\right)+\frac{\partial}{\partial y}\left(a(\Psi)\frac{\partial a(\Psi)}{\partial\Psi}\frac{\partial\phi}{\partial x}\right)\right.\left.+\phi(1-\phi)\left(\phi-\frac{1}{2}+F_{driv}+6H|\nabla\theta|\right)\right],$ (9)
@@
$\frac{\partial c_{neur}}{\partial t}=D_{c}\nabla^{2}c_{neur}+(K-k_{p75}c_{neur})\frac{\partial\phi}{\partial t}-k_{2}c_{neur},$ (12)
@@
Driving force equation. In neurons, neurotrophin concentrations are a determining factor in neuron survival. An optimal amount of neurotrophin is necessary for the survival of neurons [86]. Lower levels typically support neuron survival, while higher concentrations can negate this effect. This inverse relationship between neurotrophin concentrations and cell survival is particularly notable at higher concentrations, where an observed survival-promoting effect at lower concentrations is reversed [87]. To model the aforementioned inverse relationship between neurotrophin concentrations and cell survival considering the intricate balance of neurotrophin concentration [88], we introduce $c_{opti}$ to the driving force equation and obtain
@@
where $\frac{\alpha}{\pi}$ is a scaling coefficient, $H_{\epsilon}$ is a Heaviside step function, and $\gamma$ is the interfacial energy constant. By introducing $c_{opti}$, Eqn. 13 can effectively model the inverse relationship between $c_{neur}$ and neuron survival. When the magnitude of $c_{neur}$ exceeds $c_{opti}$, the effect of $c_{neur}$ is reversed, leading to the $\phi$ interface retractions.
@@ 4.1 Review of truncated T-spline
where $c_{i}$ is the refinement coefficients from knot insertion [89], and $N_{U^{\prime}_{i}}$ are the active children of $N_{U^{\prime}}(u)$. A bi-variate T-spline basis function, $B(u,v)=N_{U}(u)N_{V}(v)$, is the tensor product of two uni-variate functions, and its refinement is achieved by refining $N_{U}(u)$ and $N_{V}(v)$ [63].
@@
where $P_{ij}$ are the control points and $trunB_{ij}(u,v)$ are the truncated T-spline basis functions. Note that $trunB_{ij}(u,v)$ at any particular point sums up to 1 to ensure the property of partition of unity [93], which makes truncated T-splines particularly effective for accurate computational analysis with complex geometries. We refer readers to [63] for an in-depth discussion on truncated T-splines.
@@ 4.2 $\phi$ interface-based local refinements
Figure 2: Local refinements using truncated T-splines. (A) Phase field variable $\phi$ initialization on a coarse mesh. (B) Identifying elements at the $\phi$ interface. (C) Interface elements are locally refined with subdivision. (D) Detecting face-face intersections created by face extension from subdivisions. (E) Applying 1-ring of bisections to resolve these face-face intersections and maintaining the integrity of the mesh for truncated T-splines. (F) Generating locally refined T-mesh for truncated T-splines with enhanced accuracy in regions of interest. (G) Interpolating $\phi$ initialization from the coarse control mesh to the locally-refined mesh. (H) The locally refined $\phi$ initialization.
To solve the phase field model, $\phi$ initialization is needed. To apply local refinements, the model first needs to know $\phi$ initialization on a uniform coarse control mesh (Fig. 2A), refine elements at $\phi$ interface, and then obtain $\phi$ initialization on the locally refined mesh. Beginning with a uniform coarse control mesh, we can calculate $\phi$ initialization based on the initial cell center and cell radius $r_{0}$. Then, by selecting elements along $\phi$ interface (Fig. 2B), we can refine and enhance the effective mesh resolution at these critical areas of interest (Fig. 2C) while keeping the mesh coarse in areas where $\phi$ values are stable at either 0 or 1. This targeted refinement approach ensures efficient use of computational resources. Because a valid T-mesh in truncated T-splines must be strongly balanced and free of face-face intersections to satisfy linear independence [63, 90], we need to resolve these face-face intersections in the automated local refinement process (Fig. 2D). To address this issue, we apply bisections to 1-ring of elements surrounding the selected interface elements (Fig. 2E). This approach effectively resolves face-face intersections and allows the model to refine the mesh only where needed. Thus, we can obtain a locally-refined T-mesh for truncated T-splines tailored explicitly to the given $\phi$ value (Fig. 2F). Subsequently, $\phi$ values are interpolated and transferred onto this newly refined mesh using the KD tree-based interpolation method [94], providing accurate initialization for the computation of the evolving neuron outgrowth $\phi$ (Fig. 2G). Starting with an existing control point, a search using the KD tree data structure finds the nearest new control point. If a matching control point is found, the value at the input control point is directly used. If the new control point is situated among the existing points, the algorithm interpolates depending on their positions and connectivity in the control mesh (Fig. 2G). This conditional handling ensures values are correctly passed to the locally refined mesh. The KD tree approach effectively utilizes spatial indexing and positions, preserving the initialization accuracy while minimizing unnecessary computations.
@@ 5 Dynamic domain expansion
The model incorporates a dynamic domain expansion algorithm (Algorithm 1) to dynamically adjust the control mesh in response to neurite approaching the domain boundary (double-sided arrows in Fig. 3A-D). The algorithm takes neuron growth $\phi$ and control mesh as the input and evaluates each element adjacent to the domain boundary, searching for any nonzero values of the $\phi$. If nonzero $\phi$ is detected near the boundary, the algorithm flags the corresponding edge for expansion using $expFlag$. Once boundaries are flagged, the algorithm modifies the domain size based on the predetermined expansion size (default $3\times\Delta x$, where $\Delta x$ is coarse element size). It generates an enlarged control mesh (dashed zone in Figure 3A-D) and translates the entire control mesh according to the specified direction of expansion to maintain the position of the neuron $\phi$. This directional domain expansion approach works in tandem with the truncated T-splines generation, discussed in Section 4, to ensure that the computational domain adapts to accommodate growing neurites with the least amount of elements, maintaining analysis accuracy and efficiency. The transition to the expanded mesh requires an interpolation operation that passes values from the original control points to the new control points (Fig. 2G). The interpolation process for a dynamically expanding mesh in the NDDs simulation leverages the KD tree algorithm to optimize spatial queries of control points $cpts$, significantly accelerating the search for the nearest $cpts$ [94]. Starting with new control points $cpts^{\prime}$, each $cpts^{\prime}$ undergoes a rounding process to identify whether it is located on expanded elements. This procedure significantly speeds up the algorithm by skipping elements that are not expanded, and reduces the total number of DOFs while preserving the analysis accuracy.